# Countering adversarial evasion in regression analysis

### v1

@@ Abstract
Adversarial machine learning challenges the assumption that the underlying distribution remains consistent throughout the training and implementation of a prediction model. In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models, such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data. Game theoretic models have been shown to be effective at modelling these scenarios and hence training resilient predictors against such adversaries. Recent advancements in the use of pessimistic bilevel optimsiation which remove assumptions about the convexity and uniqueness of the adversary’s optimal strategy have proved to be particularly effective at mitigating threats to classifiers due to its ability to capture the antagonistic nature of the adversary. However, this formulation has not yet been adapted to regression scenarios. This article serves to propose a pessimistic bilevel optimisation program for regression scenarios which makes no assumptions on the convexity or uniqueness of the adversary’s solutions.
@@ 1 Introduction
Adversarial machine learning considers the exploitable vulnerabilities of machine learning models and the strategies needed to counter or mitigate such threats [32]. By considering these vulnerabilities during the development stage of our machine learning models, we can work to build resilient methods [9, 11] such as protection from credit card fraud [35] or finding the optimal placement of air defence systems [20]. In particular, we consider the model’s sensitivity to changes in the distribution of the data. The way the adversary influences the distribution can fall under numerous categories, see [21] for a helpful taxonomy that categorises these attacks. We focus on the specific case of exploratory attacks, which consider the scenarios where adversaries attempt to modify their data to evade detection by a classifier. Such attacks might occur in security scenarios such as malware detection [3] and network intrusion traffic [31]. In a similar vein, and more recently, vulnerabilities in deep neural networks (DNN) are being discovered, particularly in the field of computer vision and image classification; small perturbations in the data can lead to incorrect classifications by the DNN [33, 19]. These vulnerabilities raise concerns about the robustness of the machine learning technology that is being adopted and, in some cases, in how safe relying on their predictions could be in high-risk scenarios such as autonomous driving [15] and medical diagnosis [16]. By modelling the adversary’s behaviour and anticipating these attacks, we can train classifiers that are resilient to such changes in the distribution before they occur.
@@
Game theory provides an effective and popular technique to model adversarial scenarios. These games see one player, the learner, attempt to train a prediction model while another player, the adversary, modifies their data in an attempt to influence particular outcomes from the learner’s predictor. See, for example [10] for early work in modelling adversarial learning. The precise structure of such a model then depends on a number of factors, such as whether the attack occurs either at training time [27], or implementation time [5]. Further to this, while some games assume the players act simultaneously [4, 29], others allow for sequential play where either the adversary acts first [25, 26, 7, 24] or the learner acts first [21, 5]. We focus our attention on attacks made at implementation time where the adversary seeks to influence an already established prediction model, commonly known as evasion attacks in the literature [21]. Sequential evasion games naturally take on the form of a bilevel optimisation program. In this formulation, the learner, taking the role of the leader in the upper-level, train a prediction function while anticipating how the adversary, taking the role of the follower in the lower-level, will react. The solution to this program then results in a resilient prediction function. Pessimistic bilevel optimisation in particular has proved particularly effective. When multiple optimal solutions are available to the adversary, the pessimistic model, unlike its optimistic variant, assumes no cooperation between the adversary and the learner. This formulation better captures the antagonistic nature of adversarial evasion scenarios [2, 5].
@@
In this article, we adapt the pessimistic bilevel program proposed by [37] to train resilient predictors in regression scenarios. We model scenarios where we might expect adversaries to modify their data in an attempt to influence a particular label from a prediction model. We use the motivating example of an adversary providing falsified information about a house in order to influence a higher valuation. We explore properties of the constrained bilevel model under new choices of objective functions, showing the possible existence of multiple optimal solutions and hence the benefit of a constrained pessimistic bilevel approach, which, to the best of our knowledge, has not yet been applied to regression analysis. Moreover, we take this opportunity to exploit the fact that our method does not require the use of feature maps in its solution method. Feature maps, such as PCA, as was used [5], are typically not bijective. Consequently, although a solution to the adversary can be found in the map-induced feature space, a corresponding solution in the original feature space cannot be identified. However, under our model, we can view the explicit values of the adversary’s data at their solution. These values can give us insight into the strategies played by the adversary and allows us to identify the extent to which each feature needs to be modified in order to adequately influence the prediction model. From this, we can identify which features are most vulnerable. For example, if a certain feature required only a minor perturbation, then we might consider this feature to be particularly vulnerable to adversarial influence. On the other hand, if a feature required large transformations to trick the learner, then we might consider this feature to be resilient to attacks.
@@
The contributions of this article can be summarised as follows. We adapt the bilevel model proposed in [37] to accommodate regression analysis and present a novel pessimistic bilevel optimisation program to train prediction functions that are resilient to strategic changes in the data of regression scenarios. This bilevel program makes no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution, which, to the best of our knowledge, has not previously been proposed for adversarial regression analysis. We design experiments to assess the performance of our model and showcase its ability to outperform existing methods. Finally, we demonstrate a key benefit of our model over existing approaches. Where existing approaches required the use of kernel transformations to ensure convexity, we are able to keep the lower-level variable (the adversary’s data) in its original feature space, allowing us to investigate the value of the data at the solution. From this we can infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
@@
The remainder of this paper is outlined as follows; we begin in Section 2 by presenting the constrained pessimistic bilevel model to train resilient prediction functions in regression analysis before outlining the solution method. Following this, we assess the performance of the bilevel model on two datasets in Section 3 before concluding our findings in Section 4.
@@ 2 Training resilient prediction functions
In this section we present an adaptation of the constrained pessimistic bilevel program proposed in [37] to model adversarial regression. In particular, we demonstrate how the process of adversarial training can be used to train resilient predictors under the scenario of adversaries attempting to strategically modify their data in an attempt to be assigned incorrect labels from established prediction models. Such scenarios might arise, for example, in property valuation. An adversary might lie and provide falsified information in an attempt to influence a higher valuation. Through suitable choices of objective functions, we construct a constrained pessimistic bilevel program that sees a learner, in the upper-level, train a prediction function while considering how an adversary, in the lower-level, might alter their data towards a target outcome from the prediction model. The adversary is assigned some existing data as a starting point. The constraints are then used to measure the similarity between the adversary’s solution and starting point. As demonstrated in [37], these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning. In this way, the bilevel model more realistically simulates adversarial transformations, leading to improved performance. To the best of our knowledge, a constrained pessimistic bilevel program with no assumptions on convexity or uniqueness of lower-level solution has previously been proposed for adversarial evasion scenarios in regression analysis.
Let $x\in\mathbb{R}^{q}$ be a sample (row) of data containing $p\in\mathbb{N}$ features with corresponding label $y\in\mathbb{R}$. For some weights $w\in\mathbb{R}^{q}$, we define the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features,
$\sigma(w,x):=w_{0}+w^{T}x.$ (1)
This prediction function is a common choice, for example, in linear regression. We then construct a loss function, $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, to train the prediction function and identify appropriate weights. As in linear regression, we can achieve this by minimising the squared error between the learner’s prediction of $x$ and the label, $y$,
$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w_{0}+w^{T}x-y)^{2}.$
The adversary, in the lower-level, transforms some data with the aim of having the learner’s prediction function, $\sigma$ output a desired target label. For example, consider again the example of an adversary lying about features of a housing property to influence a higher valuation than deserved. If, for example, the actual valuation price is $y\in\mathbb{R}$, the adversary might set their target, $z\in\mathbb{R}$, as $z=y+\nu$, where $\nu>0$. The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label,
@@
With the loss functions established, we now work them into objective functions that accommodate a full dataset. We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then second set of data which can be manipulated by the adversary. Let $D\in\mathbb{R}^{nq},D=(D_{1},\dots,D_{n})^{T}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\{0,1\}^{n}$ be the corresponding corresponding collection of labels. Let $X\in\mathbb{R}^{mq}$ be the set of $m\in\mathbb{N}$ instances which can be manipulated by the adversary, defined as,
@@
where each $X_{i},\;i=1,\dots,m$, is a row vector of features. Let $Y\in\{0,1\}^{m}$ be the corresponding labels of the adversary’s data. The upper level objective of the bilevel program sees the learner minimising their loss over both sets of data. Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$
Then, given a set of weights, $w$ of the learner’s prediction function, the adversary, in the lower-level seeks to modify their data such that the learner’s prediction moves towards their target labels, $Z$. Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data,
@@
Finally, we define the constraints $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ as
$g(X):=\begin{pmatrix}g_{1}(X_{1})\\ \vdots\\ g_{m}(X_{m})\end{pmatrix}$, (4)
where each constraint function, $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$, is defined as
$g(X):=\delta-d(X_{i},X^{0}_{i}),$ (5)
where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function. These constraints serve to measure and restrict the cosine similarity between the adversary’s solution and initial value to be below some threshold $\delta\in\mathbb{R}$. As outlined in [37], these constraints prevent the adversary’s data becoming nonsensical and too far removed from its intended message. This continues to be useful under the regression scenarios as we can reduce the likelihood of the adversary producing unrealistic data. For example, in reality, an adversary cannot claim that their house was built thousands of years ago when in reality it was built a mere 5 years ago. This would be immediately suspicious to anybody who checks the adversary’s data. However, smaller transformations, such as within the range of $\pm 2$ years would be far more believable. Similarly to the work in [37], we use the cosine similarity to measure the extent to which the adversary has modified their data. Let $X^{0}\in\mathbb{R}^{mq}$ be the initial value of the adversary’s data, then the constraints are then defined as follows,
$g_{i}\left(X\right)=\delta-d(X_{i},X^{0}_{i})=\delta-\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\}.$ (6)
Then, setting the initial value of the adversary’s data to some existing data in the training set, we can be confident that under these constraints, the adversary produces data that are similar in nature to some real example.
We now formulate the interactions between the learner and adversary as a bilevel optimisation problem. To begin, given a set of weights, $w$, of the learner’s prediction function, let $S(w)$ be the set of optimal data for the adversary,
$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X_{i})\leq 0\right\},$ (7)
This set contains all instances of data that will trick the learner into assigning the adversary their desired labels $Z$. For example, the falsified falsified report that gives a property a better valuation than it actually deserves. Note that the adversary’s objective function under the linear regression loss in (2) and prediction function (1) possess multiple optimal solutions. This can be directly shown using in [37, Proposition 1] since the prediction function, $\sigma$, can be expressed as a function of the linear combination $w_{0}+w^{T}x$ for some $x\in\mathbb{R}^{q}$. Simply let $\Omega(w_{0}+w^{T}x)=w_{0}+w^{T}x$ and the proof follows.
@@
Note also that its has already demonstrated in [37, Proposition 2] that the feasible region defined by these constraints is non-convex. Therefore, in the case of multiple optimal solutions to the adversary’s problem, given by the set $S(w)$, we make the pessimistic assumption that the adversary will not act cooperatively, namely, they will choose the solution which maximises the learner’s objective, while the learner seeks to minimise it. The complete bilevel program is then given by
@@
The non-convexity and existence of multiple optimal solutions to the lower-level problem motives the use of the solution method proposed in [37]. To begin, we collect the upper and lower variables, $w$ and $X$, into the block variable $z$ as below
$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq},\qquad$
and define the block variable $\zeta\in\mathbb{R}^{2b+1}$ as,
$\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
where $\beta,\hat{\beta}\in\mathbb{R}^{b}$ and $\lambda\in\mathbb{R}$ are lagrange multiples with associated Lagrangian-type functions $L^{\textup{p}}_{w},L^{\textup{p}}_{X},\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ as stated below:
$L^{\textup{p}}_{w}(z,\xi):=-F(w,X)+(\beta-\lambda\hat{\beta})^{\top}g(X)$
$L^{\textup{p}}_{X}(z,\xi):=-F(w,X)+\lambda f(w,X)+\beta^{\top}g(X)$

### v2

@@ Abstract
Adversarial machine learning challenges the assumption that the underlying distribution remains consistent throughout the training and implementation of a prediction model. In particular, adversarial evasion considers scenarios where adversaries adapt their data to influence particular outcomes from established prediction models; such scenarios arise in applications such as spam email filtering, malware detection and fake-image generation, where security methods must be actively updated to keep up with the ever-improving generation of malicious data. Game theoretic models have been shown to be effective at modelling these scenarios and hence training resilient predictors against such adversaries. In this paper, we introduce a pessimistic bilevel optimsiation model, based on Stackelberg leader-follower games, to counter adversarial evasion in regression analysis. Unlike in some existing literature, we do not make assumptions (such as lower-level convexity or uniqueness of optimal solutions) on the adversary’s optimal strategy to avoid hindering their capacity and hence capture their antagonistic nature, leading to more resilient prediction models. Further to this, through the introduction of lower-level constraints, we measure and restrict the adversary’s movement, which, unlike its unrestricted counterpart in the current literature, prevents drastic transformations that do not accurately represent reality.
Keywords: Adversarial learning, game theory, bilevel optimisation, regression
@@ 1 Introduction
Adversarial machine learning considers the exploitable vulnerabilities of machine learning models and the strategies needed to counter or mitigate such threats [24]. By considering these vulnerabilities during the development stage of our machine learning models, we can work to build resilient methods [6, 8] such as protection from credit card fraud [26] or finding the optimal placement of air defence systems [15]. In particular, we consider the model’s sensitivity to changes in the distribution of the data. The way the adversary influences the distribution can fall under numerous categories, see [16] for a helpful taxonomy that categorises these attacks. We focus on the specific case of evasion attacks, which consider the scenarios where adversaries attempt to modify their data to influence particular outcomes from prediction models. Such attacks might occur in security scenarios such as malware detection [2] and network intrusion traffic [23]. In a similar vein, and more recently, vulnerabilities in deep neural networks (DNN) are being discovered, particularly in the field of computer vision and image classification; small perturbations in the data can lead to incorrect classifications by the DNN [25, 14]. These vulnerabilities raise concerns about the robustness of the machine learning technology that is being adopted and, in some cases, in how safe relying on their predictions could be in high-risk scenarios such as autonomous driving [11] and medical diagnosis [12]. By modelling the adversary’s behaviour and anticipating these attacks, we can train classifiers that are resilient to such changes in the distribution before they occur.
@@
Game theory provides an effective and popular technique to model adversarial evasion scenarios. These games see one player, the learner, attempt to train a prediction model while another player, the adversary, modifies their data in an attempt to influence particular outcomes from the learner’s predictor. See, for example [7] for early work in modelling adversarial machine learning. The precise structure of such a model then depends on a number of factors, such as whether the attack occurs either at training time [21], or implementation time [4]. Further to this, while some games assume the players act simultaneously [3, 22], others allow for sequential play where either the adversary acts first [19, 20, 5, 18] or the learner acts first [16, 4]. We focus our attention on attacks made at implementation time, where the adversary seeks to influence an already established prediction model. Sequential evasion games naturally take on the form of a bilevel optimisation program. In this formulation, the learner, taking the role of the leader in the upper-level, train a prediction function while anticipating how the adversary, taking the role of the follower in the lower-level, will react. The solution to this program then results in a resilient prediction function. Pessimistic bilevel optimisation in particular has proved particularly effective. When multiple optimal solutions are available to the adversary, the pessimistic model, unlike its optimistic variant, assumes no cooperation between the adversary and the learner. This formulation better captures the antagonistic nature of adversarial evasion scenarios [1, 27, 4].
A pessimistic bilevel formulation for evasion attacks was initially proposed in [4] to suit classification tasks, which was later extended in [1, 27] by relaxing assumptions on convexity and uniqueness of solutions, again for classification problems. In this article, we present a novel pessimistic bilevel program to train resilient predictors against evasion attacks in regression problems. We consider scenarios where we expect adversaries to modify their data in an attempt to influence a particular label from a prediction model. For example, consider a real estate surveyor who uses a prediction model to evaluate the selling price of houses. A homeowner might provide falsified information about their house, such as it’s size or distance from public services, in order to influence a higher valuation from the surveyor. Consider a simple example of two houses, $H_{1}$ and $H_{2}$. These houses are identical in every aspect except for their location. While $H_{1}$ is located just 11km from the nearest train station, $H_{2}$ is located further away at 44km. Consequently, while $H_{1}$ is evaluated at £500500k, $H_{2}$ would receive a considerably lower evaluation due to its inaccessibility to public transport. To maximise the selling price, the owner of $H_{2}$ might lie about the distance of their house as closer to the train station than it actually is. Decisions like this are simulated by the adversary in the lower-level of our bilevel program. Then, the upper-level trains the surveyor’s prediction model while considering these adversarial influences and hence trains a more resilient prediction model. Within our model, we make no assumptions about the convexity of the lower-level problem or the uniqueness of its solution, retaining the benefit provided by the pessimistic formulation.
Further to this, we introduce constraints on the lower-level optimisation problem which restrict the extent to which the adversary can modify their data. Consider again our simple housing market of houses $H_{1}$ and $H_{2}$. We established that the owner of $H_{2}$ can lie about the distance of their house from the nearest train station in order to gain a higher evaluation form the surveyor. Without restrictions on their movement, the owner of $H_{2}$ would simply state that their house is next to the train station, since a smaller distance leads to a higher evaluation. However, it is incredibility easy to spot this lie due to the large discrepancy between the true distance and the distance stated by the owner. A smaller change in distance, on the other hand, could enable the owner to receive a higher evaluation price while remaining plausibly realistic enough that the surveyor does not notice. Clearly, an adversary which takes smaller movements better reflects the strategies played by adversaries in the real world. However, without restrictions on their movement, the adversary will construct data which, while optimising their objective function, is not plausible or possibly nonsensical. The constraints we introduce here ensure that the similarity between the modified data and its original value is above some pre-defined threshold. To the best of our knowledge, existing game-theoretic approaches to adversarial regression analysis, see, for example, [28, 29, 30], do not propose a pessimistic bilevel program with lower-level constraints on the adversary’s movements that makes no assumptions on convexity or uniqueness of solutions.
@@
Moreover, where existing pessimistic approaches rely on feature maps, such as principal component analysis (PCA), or contextual embeddings to represent the adversary’s data, the adversary is allowed to manipulate their data in its original feature space. Consequently, we can view the explicit values of the adversary’s data at their solution. These values can give us insight into the strategies played by the adversary and allows us to identify the extent to which each feature needs to be modified in order to adequately influence the prediction model. From this, we can identify which features are most vulnerable. For example, if a certain feature required only a minor perturbation, then we might consider it to be particularly vulnerable to adversarial influence. On the other hand, if a feature required large transformations to trick the learner, then we might consider it to be considerably more resilient to attacks.
@@
The contributions of this article can be summarised as follows. We present a novel pessimistic bilevel optimisation program to train regression prediction functions that are resilient to strategic evasion attempts. This bilevel program, unlike existing approaches to adversarial regression, restricts the adversary’s movement through lower-level constraints, while also making no assumptions on the convexity of the lower-level problem, or the uniqueness of its solution. We design experiments to assess the performance of our model and showcase its ability to outperform existing methods. Finally, we demonstrate a key benefit of our model over existing pessimistic approaches. Where existing approaches transform the adversary’s data before analysis, removing the ability to analyse the adversary’s data at their solution, we keep it in its original feature space. This grants us the ability to infer which features require the greatest transformation in order to effectively influence the learner’s prediction function.
@@
The remainder of this article is outlined as follows; we begin in Section 2 by presenting the pessimistic bilevel model with lower-level constraints to train resilient prediction functions in regression analysis before outlining the solution method. Following this, we assess the performance of the bilevel model on two datasets in Section 3 before concluding our findings in Section 4.
@@ 2 Training resilient prediction functions
In this section we present the pessimistic bilevel program with lower-level constraints to model adversarial regression. In particular, we demonstrate how the process of adversarial training can be used to train resilient predictors under the scenario of adversaries attempting to strategically modify their data in an attempt to be assigned incorrect labels from established prediction models. Such scenarios might arise, for example, in property valuation, as outlined in the introduction. Through suitable choices of objective functions, we construct a constrained pessimistic bilevel program that sees a learner, in the upper-level, train a prediction function while considering how an adversary, in the lower-level, might alter their data towards a target outcome from the prediction model. The adversary is assigned some existing data as a starting point and the distance that the adversary diverges from this point is measured and restricted by lower-level constraints. We demonstrate, through a motivating example, how these constraints ensure that the adversary’s data remain plausibly realistic and prevent instances of data losing their intended meaning. In this way, the bilevel model more realistically simulates adversarial transformations, leading to improved performance. To the best of our knowledge, a pessimistic bilevel program with constrained adversarial movement and with no assumptions on convexity or uniqueness of lower-level solution has not previously been proposed for adversarial evasion scenarios in regression analysis.
Let $x\in\mathbb{R}^{q}$ be a sample (row) of data containing $p\in\mathbb{N}$ features with corresponding label $y\in\mathbb{R}$. For some weights $w\in\mathbb{R}^{q}$, the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, as the linear weighted sum of the features:
$\sigma(w,x):=w^{T}x.$ (1)
This is a common prediction function in linear regression. We then construct a loss function, $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$, to train the prediction function and identify appropriate weights. As in linear regression, we can achieve this by minimising the squared error between the learner’s prediction of $x$ and the label, $y$,
$\mathcal{L}(\sigma(w,x),y):=(\sigma(w,x)-y)^{2}=(w^{T}x-y)^{2}.$ (2)
The adversary, in the lower-level, transforms some data with the aim of having the learner’s prediction function, $\sigma$ output a desired target label. For example, consider again the example of an adversary lying about features of a housing property to influence a higher valuation than deserved. If, for example, the actual valuation price is $y\in\mathbb{R}$, the adversary might set their target, $z\in\mathbb{R}$, as $z=y+\nu$, where $\nu>0$. The lower-level (adversary’s) loss function is then defined as the squared distance between the learner’s prediction and the target label:
@@
With the loss functions established, we now work them into objective functions that accommodate a full dataset. We divide the training data into two sets, a static set as would typically be used in the training of a machine learning predictor, and then a second set of data which can be manipulated by the adversary. Let $D\in\mathbb{R}^{n\times q}$ be the static set of $n\in\mathbb{N}$ instances of data where each $D_{i}\in\mathbb{R}^{q},\;i=1,\dots,n$, is a row vector containing the values of $q\in\mathbb{N}$ features, and let $\gamma\in\mathbb{R}^{n}$ be the corresponding corresponding collection of labels. Let $X\in\mathbb{R}^{m\times q}$ be the set of $m\in\mathbb{N}$ instances of the same $q$ features which can be manipulated by the adversary with corresponding labels $Y\in\{0,1\}^{m}$. For convenience, we collapse the adversary’s data into a single column vector in order to reduce its dimensionality. Therefore, the adversary’s data, $X\in\mathbb{R}^{mq}$ is refined as
@@
where each $X_{i},\;i=1,\dots,m$, is a row vector of features. For consistency, we also redefine the static data in the same way, $D\in\mathbb{R}^{nq},D:=(D_{1},\dots,D_{n})^{T}$.
The upper level objective of the bilevel program sees the learner minimising their loss over both sets of data. Note that the learner considers the true labels of the adversary’s data, where $F:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ is the upper-level objective defined by
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i}).$ (4)
Then, given a set of weights, $w$ of the learner’s prediction function, the adversary, in the lower-level seeks to modify their data such that the learner’s prediction moves towards their target labels, $Z$. Let $f:\mathbb{R}^{p}\times\mathbb{R}^{mp}\rightarrow\mathbb{R}$ be the adversary’s objective, which is defined as the sum of the adversary’s loss over their data:
@@
As we have highlighted through a motivating real estate example in Section 1, an adversary who takes smaller movements better reflects the strategies played by adversaries in the real world. However, the unconstrained objective functions do not take this under consideration. The optimal adversary under the objective function in (5) will overestimate the abilities of the adversary, leading unrealistic data transformations. In order to construct a realistic adversary within the bilevel program, we introduce a set of lower-level constraints on the adversary’s optimization problem which measure the similarity between the true value of the adversary’s data and the value seen by the learner. Let $X^{0}$ be the true (original) value of the adversary’s data, and let $g:\mathbb{R}^{mq}\rightarrow\mathbb{R}^{m}$ be the vector of constraint functions
$g(X):=\begin{pmatrix}g_{1}(X)\\ \vdots\\ g_{m}(X)\end{pmatrix}$. (6)
We define each constraint function component $g_{i}:\mathbb{R}^{q}\rightarrow\mathbb{R}$, for each instance of the adversary’s data $i=1,\ldots,m$, by
$g_{i}(X):=\delta-d(X_{i},X^{0}_{i}),$ (7)
where $d:\mathbb{R}^{q}\times\mathbb{R}^{q}\rightarrow\mathbb{R}$ is some similarity function and $\delta$ is the minimum required similarity threshold. The constraint $g(X)\leq 0$ then restrict the adversary to produce data whose similarity is greater than $\delta$. The work in [27] demonstrated the effectiveness of the cosine similarity to measure the extent to which the adversary has modified their data and so we use the same here,
$d(X_{i},X^{0}_{i}):=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i=\{1,\dots,m\}.$ (8)
Combining the lower-level objective 5 with the constraints in 8 gives the adversary’s optimisation problem, the solution to which provides data that trick the learner into giving the best possible label assignment while remaining plausibly realistic. Given a set of weights of the learner’s prediction function, $w$, we defined $S(w)$ as this set of optimal data, given by
$S(w):=\operatorname*{argmin}_{X\in\mathbb{R}^{mq}}\left\{f(w,X)\mid g(X)\leq 0\right\}.$ (9)
@@
Note that the adversary’s objective function under the linear regression loss in (3) and prediction function (1) possess multiple optimal solutions. This can be directly shown using [27, Proposition 1] since the prediction function, $\sigma(w,x)=w^{T}x$, is a linear combination of $w$ and $x$. Note also that it has already been demonstrated in [27, Proposition 2] that the feasible region defined by the constraints in (8) is non-convex. Therefore, in the case of multiple optimal solutions to the adversary’s problem, given by the set $S(w)$, we make the pessimistic assumption that the adversary will not act cooperatively, namely, they will choose the solution which maximises the learner’s objective, while the learner seeks to minimise it. The complete bilevel program is then given by
@@
A solution to the problem described by (9) - (10) will comprise the weights of a prediction that has accounted for adversarial manipulation during its training process. Consequently, the prediction function will be more resilient to evasion attacks made after implementation. We say that a set of weights $\bar{w}$ is a local optimal solution for problem (9) - (10) if there is a neighbourhood $W$ of the point such that
$\varphi_{p}(\bar{w})\leq\varphi_{p}(w)\;\,\mbox{ for all }\;w\in W,$ (11)
where the $\varphi_{p}$ denotes the following two-level value function (concept first introduced and studied in [10]):
$\varphi_{\textup{p}}(w):=\max\{F(w,X)\,|\,X\in S(w)\}.$
[…]
$\nabla_{w}F(w,X)$ $=0,$ (12a)
$\nabla_{X}F(w,X)-\lambda\nabla_{X}f(w,X)-\nabla g(X)^{\top}\beta$ $=0,$ (12b)
$\nabla_{X}f(w,X)+\nabla g(X)^{\top}\hat{\beta}$ $=0,$ (12c)
$\hat{\beta}\geq 0,\quad g(X)\leq 0,\quad\hat{\beta}^{\top}g(X)$ $=0,$ (12d)
$\lambda\geq 0,\;\,\beta\geq 0,\quad g(X)\leq 0,\quad\beta^{\top}g(X)$ $=0,$ (12e)
for some data $X$ belonging to the adversary. We can solve this system by first transforming it into a system of equations. To simplify the notation, we introduce the block variables
$z:=\begin{bmatrix}w\\ X\end{bmatrix}\in\mathbb{R}^{q+mq}\;\mbox{ and }\;\xi:=\begin{bmatrix}\beta\\ \hat{\beta}\\ \lambda\end{bmatrix}\in\mathbb{R}^{2b+1},$
as well as the block functions
$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix}\;\mbox{ and }\;\mathcal{H}(z,\xi):=\begin{bmatrix}\nabla_{w}F(w,X)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$,
where $L^{\textup{p}}_{X},\,\ell^{\textup{p}}\colon(\mathbb{R}^{q}\times\mathbb{R}^{mq})\times\mathbb{R}^{2b+1}\to\mathbb{R}$ are upper and lower-level Lagrangian type functions:
$L^{\textup{p}}_{X}(z,\xi):=F(w,X)-\lambda f(w,X)-\beta^{\top}g(X),$